OpenStudy (anonymous):

Find an equation of the line that satisfies the given conditions. A. Write the equation in standard form. B. Write the equation an slope intercept form. x intercept is (2,0); Slope is 6.

OpenStudy (anonymous):

Can someone help me out

OpenStudy (anonymous):

How would I write this in standard form

OpenStudy (anonymous):

A. the equation for standard form is ax+by=c (plug what your given in including the slope) B. slope intercept form is y=mx+b (do the same and plug in what you have)

OpenStudy (anonymous):

6x+2=0. is this right

OpenStudy (anonymous):

I am having trouble with this problem

OpenStudy (ahsome):

A linear equation is commonly written in slope intercept form form,\(y=mx+b\) Where: \(y\) and \(x\) are two variables that change \(m\) Is the Slope \(b\) Is the constant, which is equal to the \(y-intercept\) If the x intercept is \(2\), then you know that the \(y\) value is 0 And since you know that \(m = 6\), now all that's left to do is to find \(b\). \(\text{Use Slope-Intercept Form}\)\[y=mx+b\]Sun in your values\[y=mx+b\]\[0=6\times2+b\]\[0=12+b\]\[-b=12\]\[b=-12\] Hence, for slope intercept form, \(y=6x-12\) To turn this form to standard form, simply move the \(6x\) to the LHS (Left Hand Side) \[y=6x-12\]\[y-6x=-12\]\[-6x+y=-12\] Thats it :). Tell me if that doesn't make sense, @Bigtim10

OpenStudy (anonymous):

Yes it make since now that I see it written

OpenStudy (anonymous):

Thanks for your help